# Linear viscoelastic rheological FrBD models

### v2

@@ 4.2 Feedback control of a robotic arm
From Proposition 2 in [11], it follows that the operator $\mathrm{\Sigma}_{1}\mathrel{\mathop{\ordinarycolon}}-\tilde{F}\mapsto s$ defined by (29a) is output-strictly-passive, that is, there exists $\beta_{1}\in\mathbb{R}_{>0}$ such that
$-\int_{0}^{t}\tilde{F}\bigl(t^{\prime}\bigr)s\bigl(t^{\prime}\bigr)\operatorname{d\!}t^{\prime}\geq\beta_{1}\int_{0}^{t}\mathinner{\!\left\lvert s\bigl(t^{\prime}\bigr)\right\rvert}^{2}\operatorname{d\!}t^{\prime},\quad t\in\mathbb{R}_{\geq 0}.$ (30)
Moreover, by Lemma 3.4, the operator $\mathrm{\Sigma}_{2}\mathrel{\mathop{\ordinarycolon}}s\mapsto\tilde{F}$ defined by (29b), (29c), and $\tilde{F}=rp\tilde{f}$ is passive. In particular, there exists $\beta_{2}\in\mathbb{R}_{\geq 0}$ such that
$\int_{0}^{t}\tilde{F}\bigl(t^{\prime}\bigr)s\bigl(t^{\prime}\bigr)\operatorname{d\!}t^{\prime}\geq-\beta_{2}.$ (31)

### v3

@@ 4.2 Feedback control of a robotic arm
From Proposition 2 in [11], it follows that the operator $\mathrm{\Sigma}_{1}\mathrel{\mathop{\ordinarycolon}}-\tilde{F}\mapsto s$ defined by (29a) is output-strictly-passive, that is, there exists $\beta_{0}\in\mathbb{R}_{\geq 0}$ and $\beta_{1}\in\mathbb{R}_{>0}$ such that
$-\int_{0}^{t}\tilde{F}\bigl(t^{\prime}\bigr)s\bigl(t^{\prime}\bigr)\operatorname{d\!}t^{\prime}\geq-\beta_{0}+\beta_{1}\int_{0}^{t}\mathinner{\!\left\lvert s\bigl(t^{\prime}\bigr)\right\rvert}^{2}\operatorname{d\!}t^{\prime},\quad t\in\mathbb{R}_{\geq 0}.$ (30)
Moreover, by Lemma 3.4, the operator $\mathrm{\Sigma}_{2}\mathrel{\mathop{\ordinarycolon}}s\mapsto\tilde{F}$ defined by (29b), (29c), and $\tilde{F}=rp\tilde{f}$ is passive. In particular, there exist $\beta_{2}\in\mathbb{R}_{\geq 0}$ such that
$\int_{0}^{t}\tilde{F}\bigl(t^{\prime}\bigr)s\bigl(t^{\prime}\bigr)\operatorname{d\!}t^{\prime}\geq-\beta_{2},\quad t\in\mathbb{R}_{\geq 0}.$ (31)